# Language Models can Self-Improve at State-Value Estimation for Better Search

### v1

@@ 7 Conclusion
We propose self-taught lookahead (STL) as an efficient method to improve the value model employed during search. This efficiency primarily takes the form of reward and human demonstration-free improvement, where models instead learn from state-transition dynamics. However, we also show that because STL allows us to perform self-improvement on small models that can be deployed with less exhaustive search algorithms, search with STL models yields significant reductions in both cost and environmental usage. Therefore, the STL framework could help enable the more realistic training and deployment of agent systems.
Impact Statement
Our STL method enables LLMs to self-improve at search by capturing the mechanics of traditional RL algorithms in natural language. While we leverage this technique for better state-value estimation, it can likely apply to other settings. Additionally, we show in §5, that STL requires significantly less compute overhead than similarly performing methods, thus reducing the energy consumption required for search during inference, helping to improve the sustainability of agent deployment. Finally, having models self-improve without human supervision may enable agents to take unintended and potentially harmful actions. While investigating these harms is out of the scope of this work, we encourage future research in this area.
@@ Appendix A Self-Taught Lookahead Algorithm
The STL algorithm is presented in full in Algorithm 1. For information about the task_specific_filter, see Appendix B.3 and Appendix C.3.
@@
Appendix B Self-Taught Lookahead on WebShop
In this section, we outline information about WebShop (Yao et al., 2022a) and the implementation details of running STL on the benchmark.
B.1 WebShop Task
@@
    Search actions allow the user to search for a particular item with a natural language query e.g. search[easy to use medium color face kit less than 40 dollars]. This action can only be taken on the search page which is also the initial / home page of the WebShop interface.
  • click[button]: Click actions are discrete actions but can take many forms which we enumerate below:
@@
    • click[attribute]: to toggle on an attribute or option on the product page of an item e.g. click[small]
    • click[Buy Now]: to buy the selected item - this is a terminal action that yields the ground truth reward. This action is not allowed to be taken in STL search but is allowed in other search, RL, and prompting methods (Yao et al., 2022a; Shinn et al., 2024; Zhou et al., 2024).
    • Other navigation buttons: other navigation buttons include click[Back to Search], click[<Prev], click[Next>], click[Description], click[Features]. To simplify trajectories, we generally restrict the ability for models to take these actions in all settings following (Yao et al., 2022a).
WebShop provides a textual representation of webpages in simple mode. An example of this representation for search results are shown in Figure 6.
@@
B.2 Prompts
The prompt used to generate actions in WebShop is presented in Figure 7. Notice that we do not use think actions part of the classical ReACT framework (Yao et al., 2022b) like Yao et al. (2022a) or Zhou et al. (2024) because evaluating the value of these actions is difficult as they have no observation. Instead, we prompt the policy model to provide a rationale while generating possible actions. Also, note that we prompt the policy multiple times adding to the list of actions that are not allowed and removing from the list of actions that are allowed. This change to a “selection” policy enables action diversity which we find is otherwise low even with prompting the policy at high temperature. This change also likely explains why there is no statistically significant difference between using a gpt-3.5-turbo and gpt-4o policy in §4.1.
Likewise, the prompt used to evaluate states is presented in Figure 8. Note that this evaluation prompt is only used to prompt base models, STL value models are only prompted with the current trajectory. We note that the Likert scale used was crucial to obtaining consistent value outputs on which we could perform STL. For all value estimates (base model or fine-tuned), we prompt the value model 5 times and use the average score as the state value estimate. During the data generation phase, since we need a single rationale to fine-tune on which to construct the action-outcome rationale, we choose the rationale corresponding to the median of the 5 scores.
WebShop Generation Prompt You are a web agent, select the best next action for the search to fulfill the task. Example tasks are shown below. Provide a rationale for your selection BEFORE you provide the action. NOTE: You can only select actions that are provided in the Possible Actions list. You MAY NOT select actions in the Not Allowed list. NOTE: You must output BOTH a rationale and an action. NOTE: Do not select any of the following actions: ’Back to Search’, ’Next >’, ’< Prev’, ’Attributes’, ’Description’, ’Features’, ’Reviews’, even if they are available on the page. Example Tasks: {few shot examples} —————————————————————————————————————————– New Task: {task} Actions Not Allowed: {not_allowed_actions} Possible Next Actions (REMINDER: You can only select actions from this list.): {possible_actions} REMINDER: Do not select any of the following actions: ’Back to Search’, ’Next >’, ’< Prev’, ’Attributes’, ’Description’, ’Features’, ’Reviews’, even if they are available on the page.
@@ B.2 Prompts
WebShop Evaluation Prompt Given an item to purchase and a trajectory that aims to buy an item that exactly matches the specification, analyze how well the last action and observation align with the task. Provide a reflection that concludes with. ”Thus the correctness score is s”, where s is either 1, 2, 4, 6, 8, or 10. Use the following scale for scoring: 1: The last action and observed state is entirely irrelevant to the task or captures a purchase of an item that is completely unrelated to the specifications. 2: The last action and observed state captures a step with a low likelihood of leading to purchasing the correct item. 4: The last action and observed state captures a step with a moderate likelihood of leading to purchasing the correct item. 6: The last action and observed state captures a step with a high likelihood of leading to purchasing the correct item. 8: The last action and observed state captures a step with a very high likelihood of leading to purchasing the correct item. 10: The last action and observed state captures a step that will certainly lead to purchasing the correct item. Keep reflections short (<100 words). Follow the format of the rationale from the below example task. NOTE: the observation from clicking on the item will be the item’s product detail page. For instance, click[B078GWRC1J] will show the product detail page for the item with code B078GWRC1J which will include the item’s name (e.g. Bright Citrus Deodorant by Earth Mama), price (\$10.99), and other relevant details as well as options. NOTE: Assume none of the attributes on the product page are selected only provide the reflection for the last action. Example Tasks: {few shot examples} —————————————————————————————————————————– New Task: Respond with the reflection for the last observation of the new task ONLY. As a remind,er the last action and observation is as follows: {last_action} Your response should start with ”Reflection:” and end with ”Thus the correctness score is …”.
@@
Figure 8: Evaluation prompt for WebShop. This prompt was only used to prompt base models.
B.3 Implementing STL
@@
Unlike with math reasoning tasks, with web tasks, the position of an action in a trajectory may influence its value. For instance, selecting a certain item I from search results early in the trajectory should have a higher value than selecting I a second time in the same trajectory. To account for this difference, we train value models at each position (depth) in the trajectory. Specifically, we limit trajectories to five steps and train four values models depths 1 to 4 only allowing a terminating Buy action on the final step.
@@
We also filter out malformed rationales from the training data. Specifically, we remove rationales that do not provide the proper format e.g. it does not exactly contain scaffolding like “Thus the correctness score is”.
In total, including scaffolding, the generated self-improvement dataset consists of 1210448 tokens.
B.4 Difficulty in Performing STL for multiple iterations.
@@
Appendix C Math Tasks
C.1 Game-of-24 Task
@@
As introduced by Yao et al. (2024), the Game-of-24 is a mathematical reasoning task that involves combining four numbers e.g. “2 3 4 5” together with mathematical operations i.e. $+$, --, ///, $\times$ in order to obtain 24. An action in this task consists of simply applying a mathematical operation to combine two numbers e.g. $2+3=5$, the resulting state from the operation is the set of remaining numbers e.g “5 4 5”.
C.2 Prompts
@@
The prompt used to generate actions for the Game-of-24 task is presented in Figure 9. Likewise, the prompt used to evaluate states is presented in Figure 10. Like WebShop, this evaluation prompt is only used to prompt base models, STL value models are only prompted with the current trajectory. However, unlike WebShop, values are not real numbers between 1 and 10, but rather 0.001, 1, and 20 corresponding to the labels of impossible, likely, sure that the remaining numbers can be combined to reach 24. Note these values were used by the Yao et al. (2024) in the original Tree-of-Thoughts paper, but are ad-hoc and are used purely as labels. For all value estimates (base model or fine-tuned), we prompt the value model 3 times and use the median score as the state value estimate. During the data generation phase, since we need a single rationale to fine-tune on which to construct the action-outcome rationale, we choose the rationale corresponding to the median of the 3 scores.
C.3 Implementing STL
@@
In total, including scaffolding, the generated self-improvement dataset consists of 28708 tokens.
Game-of-24 Generation Prompt Use numbers and basic arithmetic operations (+ - * /) to obtain 2In eachach step, you are only allowed to choose two of the remaining numbers to obtain a new number. Follow the example format exactly. {few shot examples} —————————————————————————————————————————– {input}
Figure 9: Generation prompt for the Game-of-24 policy.
Game-of-24 Evaluation Prompt Evaluate if given numbers can reach 24 (sure/likely/impossible) Follow the example format exactly. Only evaluate the last example. few shot examples —————————————————————————————————————————– {input}
Figure 10: Evaluation prompt for Game-of-24. This prompt was only used to prompt base models.
Appendix D Model Fine-tuning and Serving
@@
Table 4: Hyperparameters during STL training.
Fine-tuning the value model for STL is carried out on a single NVIDIA A40 GPU. We use LoRA finetuning (Hu et al., 2021) and use models provided by unsloth888unsloth.ai. The hyperparameters used are in Table 4. We fine-tuned Game-of-24 value models for 10 epochs and WebShop value models for 20 epochs due to the differences in difficulty for models to learn the format of the action and state representations.
Additionally, we serve base and fine-tuned models using vLLM 999docs.vllm.ai/en/latest/ (Kwon et al., 2023) for efficient value estimation of new states during search.
Appendix E Significance Testing
In §4.1, we use the paired bootstrap test to test the statistical significance of our experimental results. Following Berg-Kirkpatrick et al. (2012), we set $b=10^{6}$. We also run the significance test twice: once for score (average reward) and a separate time for success rate.

### v2

@@ 7 Conclusion
We propose STL as an efficient method to improve the value model employed during search. This efficiency primarily stems from STL’s design for information-scarce settings, where models learn from state-transition dynamics. Additionally, because STL enables self-improvement on small models deployable with less exhaustive search, it yields significant reductions in both computational cost and environmental usage. Therefore, the STL framework could help enable the more realistic learning and deployment of agent systems.
@@ Appendix A Self-Taught Lookahead Algorithm
The STL algorithm is presented in full in Algorithm 1. For information about the task_specific_filter, see Appendix D.3 and Appendix F.4.
@@
Appendix B Broader Impacts
Our STL method enables LLMs to self-improve at search by capturing the mechanics of traditional RL algorithms in natural language. While we leverage this technique for better state-value estimation, it can likely be applied to other problems. Additionally, we show in §5, that STL requires significantly less compute overhead than similarly performing methods, thus reducing the energy consumption required for search during inference, helping to improve the sustainability of agent deployment.
As with all self-improvement work [43, 61], having models self-improve without human supervision may enable agents to learn how to take actions towards task completion that are not well-aligned with human values or preferences. To prevent misuse, we only release code to reproduce experiments in this paper and not to improve general-purpose agents for which these risks might be more prevalent. While investigating these harms is out of the scope of this work, we encourage future research in this area.
Appendix C Limitations
During self-improvement, STL does require task specifications like “I am looking for a queen-sized bed that is black …’ (see Figure 2) for web tasks. However, this assumption is well-founded as prior self-improvement work, such as that on LLM alignment [62, 54], also assumes tasks, or in their case, user prompts, are provided to initialize the self-improvement process. Additionally, as mentioned throughout the paper, this setting is much more reasonable than Reward and Demo Learning and Reward-Guided Inference settings, which require ground truth reward and/or human demonstrations along with task specifications.
Additionally, due to compute constraints, our experiments utilize models up to 8 billion parameters for STL. However, our results in §5 demonstrate a scaling trend in performance with respect to model size, indicating that using larger models may yield more performant agents. Given these positive results, we leave applying STL to larger models to teams with larger resource budgets.
Appendix D Self-Taught Lookahead on WebShop
In this section, we outline information about WebShop [58] and the implementation details of running STL on the benchmark.
D.1 WebShop Task
@@
    Search actions allow the user to search for a particular item with a natural language query, e.g., search[easy to use medium color face kit less than 40 dollars]. This action can only be taken on the search page, which is also the initial / home page of the WebShop interface.
  • click[button]: Click actions are discrete actions but can take many forms, which we enumerate below:
@@
    • click[attribute]: to toggle on an attribute or option on the product page of an item, e.g. click[small]
    • click[Buy Now]: to buy the selected item - this is a terminal action that yields the ground truth reward. This action is not allowed to be taken in STL search but is allowed in other search, RL, and prompting methods [58, 46, 68].
    • Other navigation buttons: other navigation buttons include click[Back to Search], click[<Prev], click[Next>], click[Description], click[Features]. To simplify trajectories, we generally restrict the ability for models to take these actions in all settings following [58].
WebShop provides a textual representation of webpages in simple mode. An example of this representation for search results is shown in Figure 7.
@@
D.2 Prompts
The prompt used to generate actions in WebShop is presented in Figure 8. Notice that we do not use think actions part of the classical ReACT framework [60] like [58] or [68] because evaluating the value of these actions is difficult as they have no observation. Instead, we prompt the policy model to provide a rationale while generating possible actions. Also, note that we prompt the policy multiple times, adding to the list of actions that are not allowed and removing from the list of actions that are allowed. This change to a “selection” policy enables action diversity, which we find is otherwise low even with prompting the policy at high temperature. This change also likely explains why there is no statistically significant difference between using a gpt-3.5-turbo and gpt-4o policy in §4.1.
Likewise, the prompt used to evaluate states is presented in Figure 9. Note that this evaluation prompt is only used to prompt base models; STL value models are only prompted with the current trajectory. We note that the Likert scale used was crucial to obtaining consistent value outputs on which we could perform STL. For all value estimates (base model or fine-tuned), we prompt the value model 5 times and use the average score as the state value estimate. During the data generation phase, since we need a single rationale to fine-tune on which to construct the action-outcome rationale, we choose the rationale corresponding to the median of the 5 scores.
WebShop Generation Prompt You are a web agent, select the best next action for the search to fulfill the task. Example tasks are shown below. Provide a rationale for your selection BEFORE you provide the action. NOTE: You can only select actions that are provided in the Possible Actions list. You MAY NOT select actions in the Not Allowed list. NOTE: You must output BOTH a rationale and an action. NOTE: Do not select any of the following actions: ’Back to Search’, ’Next >’, ’< Prev’, ’Attributes’, ’Description’, ’Features’, ’Reviews’, even if they are available on the page. Example Tasks: {few shot examples} —————————————————————————————————– New Task: {task} Actions Not Allowed: {not_allowed_actions} Possible Next Actions (REMINDER: You can only select actions from this list.): {possible_actions} REMINDER: Do not select any of the following actions: ’Back to Search’, ’Next >’, ’< Prev’, ’Attributes’, ’Description’, ’Features’, ’Reviews’, even if they are available on the page.
@@ D.2 Prompts
WebShop Value Estimation Prompt Given an item to purchase and a trajectory that aims to buy an item that exactly matches the specification, analyze how well the last action and observation align with the task. Provide a reflection that concludes with. "Thus the correctness score is s", where s is either 1, 2, 4, 6, 8, or 10. Use the following scale for scoring: 1: The last action and observed state is entirely irrelevant to the task or captures a purchase of an item that is completely unrelated to the specifications. 2: The last action and observed state captures a step with a low likelihood of leading to purchasing the correct item. 4: The last action and observed state captures a step with a moderate likelihood of leading to purchasing the correct item. 6: The last action and observed state captures a step with a high likelihood of leading to purchasing the correct item. 8: The last action and observed state captures a step with a very high likelihood of leading to purchasing the correct item. 10: The last action and observed state captures a step that will certainly lead to purchasing the correct item. Keep reflections short (<100 words). Follow the format of the rationale from the below example task. NOTE: the observation from clicking on the item will be the item’s product detail page. For instance, click[B078GWRC1J] will show the product detail page for the item with code B078GWRC1J which will include the item’s name (e.g. Bright Citrus Deodorant by Earth Mama), price (\$10.99), and other relevant details as well as options. NOTE: Assume none of the attributes on the product page are selected only provide the reflection for the last action. Example Tasks: {few shot examples} —————————————————————————————————– New Task: Respond with the reflection for the last observation of the new task ONLY. As a reminder the last action and observation is as follows: {last_action} Your response should start with "Reflection:" and end with "Thus the correctness score is …".
@@
Figure 9: Value Estimation prompt for WebShop. This prompt was only used to prompt base models.
D.3 Implementing STL
@@
With web tasks, the position of an action in a trajectory may influence its value. For instance, selecting a certain item I from search results early in the trajectory should have a higher value than selecting I a second time in the same trajectory. To account for this difference, we train value models at each position (depth) in the trajectory. Specifically, we limit trajectories to five steps and train four value models depths 1 to 4, only allowing a terminating Buy action on the final step.
@@
We also filter out malformed rationales from the training data. Specifically, we remove rationales that do not provide the proper format, e.g., it does not exactly contain scaffolding like “Thus the correctness score is”.
Additionally, we generate lookahead rollouts using 5 iterations of MCTS, which, in practice, we find is sufficient for collecting diverse training data.
D.4 Difficulty in Performing STL for multiple iterations.
@@
Appendix E Self-Taught Lookahead on HotpotQA
E.1 HotpotQA Task
The HotpotQA [57] is a multi-hop question answering benchmark where correct answers require reasoning over multiple Wikipedia entries. There are three possible actions at each step:
  • search[entry]:
    Search provides the first five sentences of the corresponding Wikipedia entry if it exists, or provides five alternative existing entries.
  • lookup[string]:
    Lookup returns the next sentence in the entry containing the specified string.
  • finish[string]:
    Finish signifies the completion of the reasoning process, where the answer is specified with the provided string.
E.2 Prompts
The prompt used to generate actions for the HotpotQA task is presented in Figure 12. Likewise, the prompt used to evaluate states is presented in Figure 11. Many of the prompting details, such as the use of a disallowed action list, are similar to the WebShop task. However, one difference is that if five unique actions are not found due to a search term not being an entry, we add the top-5 similar terms returned by the retrieval model as possible actions.
E.3 Evaluation
We use the same gpt-4o evaluation prompt as [13] which is provided in Figure 11. Note that a match is computed when either there is an exact match or a match indicated by the gpt-4o evaluator in order to prevent any mistakes by the evaluator, such as indicating a non-match even though the two answers were an exact string match.
Additionally, unlike WebShop, we limit trajectories to a depth of four, and explicitly prompt models to provide a final answer on the fourth step in the trajectory.
E.4 Implementing STL
The implementation details of STL for HotpotQA are similar to WebShop in the training of multiple models at different depths, the filtering out of malformed rationales, and the use of 5 iterations of MCTS during data generation. However, one difference is that we feed the possible next actions to the value model so that it can provide coherent lookahead simulations and evaluations.
HotpotQA Generation Prompt Solve a question answering task with interleaving Thought, Action, Observation steps. Thought can reason about the current situation, and Action can be three types: (1) Search[entity], which searches the exact entity on Wikipedia and returns the first paragraph if it exists. If not, it will return some similar entities to search. (2) Lookup[keyword], which returns the next sentence containing keyword in the current passage. (3) Finish[answer], which returns the answer and finishes the task. After each observation, provide the next Thought and next Action. NOTE: You MAY NOT select actions in the Actions Not Allowed list. You have to change the wording of the query or lookup somehow. NOTE: Keep search queries and lookups short and concise since longer queries will not return any result. For instance, instead of searching Search[eastern sector of the Colorado orogeny], search Search[Colorado orogeny] and then Lookup[eastern sector]. Here are some examples: {few shot examples} —————————————————————————————————- New Task: {task} Actions Not Allowed: {not_allowed_actions} Thought:
Figure 10: Generation prompt for the HotpotQA policy.
HotpotQA Value Estimation Prompt Given a question and a trajectory to answer the question, analyze how well the LAST ACTION in the trajectory contributes to finding the answer. Consider ONLY the last action. The trajectories are labeled by pairs of thoughts that can reason about the current situation and actions that can be of three types: (1) Search[entity], which searches the exact entity on Wikipedia and returns the first paragraph if it exists. If not, it will return some similar entities to search. (2) Lookup[keyword], which returns the next sentence containing keyword in the current passage. (3) Finish[answer], which returns the answer and finishes the task. Provide a reflection that concludes with “Thus the correctness score is s”, where s is either 1, 3, 5, 7, or 10. Use the following scale for scoring: 1: The action is completely irrelevant to answering the question or there is no relevant search result (“Could not find” is in the observation). 3: The action’s observation provides information only at the background level to answering the question. 5: The action’s observation provides information that makes a small step towards answering the question. 7: The action’s observation provides information that makes a key step towards answering the question. 10: The action’s observation provides the final piece of information needed to answer the question. Reminder: If you see “Could not find” in the observation, the correctness score is 1. Keep reflections short (< 100 words). Follow the following examples: {few shot examples} —————————————————————————————————- {input}
Figure 11: Value estimation prompt for HotpotQA. This prompt was only used to prompt base models.
HotpotQA Evaluation Prompt Based on the provided question and reference answer, please determine if the response is correct or incorrect. Begin by articulating your rationale, and conclude with a single word judgment: ‘Yes’ for correct or ‘No’ for incorrect. question: {question} reference answer: {reference} response: {response}
Figure 12: Evalaution prompt for the HotpotQA predicted answers.
Game-of-24 Generation Prompt Use numbers and basic arithmetic operations (+ - * /) to obtain 24. In each step, you are only allowed to choose two of the remaining numbers to obtain a new number. Follow the example format exactly. {few shot examples} —————————————————————————————————- {input}
Figure 13: Generation prompt for the Game-of-24 policy.
Game-of-24 Value Estimation Prompt Evaluate if the given numbers can reach 24 (sure/likely/impossible) Follow the example format exactly. Only evaluate the last example. {few shot examples} —————————————————————————————————- {input}
Figure 14: Value estimation prompt for Game-of-24. This prompt was only used to prompt base models.
Appendix F Self-Taught Lookahead on Game-of-24
F.1 Game-of-24 Task
@@
As introduced by [59], the Game-of-24 is a mathematical reasoning task that involves combining four numbers e.g. “2 3 4 5” together with mathematical operations i.e. $+$, --, ///, $\times$ in order to obtain 24. An action in this task consists of simply applying a mathematical operation to combine two numbers e.g. $2+3=5$, the resulting state from the operation is the set of remaining numbers e.g “5 4 5”.
F.2 Prompts
@@
The prompt used to generate actions for the Game-of-24 task is presented in Figure 13. Likewise, the prompt used to evaluate states is presented in Figure 14. Like WebShop, this evaluation prompt is only used to prompt base models; STL value models are only prompted with the current trajectory. However, unlike WebShop, values are not real numbers between 1 and 10, but rather 0.001, 1, and 20, corresponding to the labels of impossible, likely, and sure that the remaining numbers can be combined to reach 24. Note that these values were used by the original Tree-of-Thoughts paper [59], but are ad-hoc and are used purely as labels. For all value estimates (base model or fine-tuned), we prompt the value model 3 times and use the median score as the state value estimate. During the data generation phase, since we need a single rationale to fine-tune on which to construct the action-outcome rationale, we choose the rationale corresponding to the median of the 3 scores.
F.3 Evaluation
For all tested BFS methods, we use the same setup as the Tree-of-Thoughts paper, i.e., we select 5 of the best actions (set the branching factor to 5) at each of the 4 steps (two numbers are combined during each step).
F.4 Implementing STL
@@
Unlike with WebShop and HotpotQA, we do not train value models at each depth, due to the small state space. Additionally, we also use a branching factor of 5 during data generation.
Appendix G Costs
Here we detail how we computed costs in the efficiency analysis in §5.
G.1 Inference Costs
To compute inference cost, we use OpenAI’s 444openai.com/api/pricing for OpenAI models like gpt-3.5-turbo and gpt-4o and Groq’s 555groq.com/pricing pricing tables for closed source llama models. Note that we only use Groq for a pricing estimate, as inference and training were run on in-house GPUs. As costs may change over time, we provide the pricing figures that we used for our experiments in Table 4.
Table 4: Inference pricing used for cost analysis.
Prompt Tokens (\$ / 1000 tokens) Completion Tokens (\$ / 1000 tokens)
gpt-3.5-turbo 0.0005 0.0015
gpt-4o 0.0025 0.01
llama-3.1-8b-instruct 0.00005 0.00008
G.2 Fine-tuning Costs
We also account for the costs incurred from fine-tuning llama-3.1-8b-instruct during STL. Fine-tuning with a single A40 GPU takes 4.5 hours for the WebShop task. While it is difficult to estimate costs on our own in-house GPUs, we can estimate the cost by using VastAI’s figure of \$0.39 per A40 GPU hour  [^6] https://vast.ai/pricing/gpu/A40 when these experiments were run in April, 2025. Thus, the total training run cost roughly \$1.76. We also note that as the number of tasks seen at test time scales, the fine-tuning costs become increasingly negligible.
Appendix H Model Fine-tuning and Serving
Table 5: Hyperparameters during STL training.
@@ Appendix H Model Fine-tuning and Serving
Table 6: Effect of $\gamma$ on Performance.
$\gamma$ $0.75$ $0.80$ $0.85$ $0.90$ $0.95$ $0.99$ $1.0$
WebShop Success Rate $32.0$ $24.0$ $24.0$ $24.0$ $28.0$ $42.0$ $46.0$
Fine-tuning the value model for STL is carried out on a single NVIDIA A40 GPU. We use LoRA finetuning [23] and use models provided by unsloth777unsloth.ai. The hyperparameters used are in Table 5. We fine-tuned Game-of-24 value models for 10 epochs and WebShop value models for 20 epochs due to the differences in difficulty for models to learn the format of the action and state representations.
Additionally, we serve base and fine-tuned models using vLLM 888docs.vllm.ai/en/latest/ [31] for efficient value estimation of new states during search.
Finally, we perform a small analysis on the effect of the discount factor $\gamma$ on search performance. The results of the analysis with a gpt-3.5-turbo policy on WebShop are presented in Table 6. These results confirm the choice of $\gamma=1.0$ from prior work cited in §3.
Appendix I Significance Testing
In §4.1 and §4.2, we use the paired bootstrap test to test the statistical significance of our experimental results. Following [3], we set $b=10^{6}$. For WebShop, we run the significance test twice: once for score (average reward) and a separate time for success rate.